OpenStudy (ucreategames):

geoff has $5000 to invest. he wants to earn $568 in interest in one year. he will invest part of the money at 12% and the other part at 10%. how will he invest to earn $568.

OpenStudy (lilg132):

5000 = x + y 5000 - x = y 0.1x + 0.12y = 568 0.1x + 0.12(5000 - x) = 568 0.1x + 600 - 0.12x = 568 600 - 568 = 0.12x - 0.1x 32 = 0.02x 32 / 0.02 = 0.02x / 0.02 x = 1600 0.1x + 0.12y = 568 0.1(1600) + 0.12y = 568 160 + 0.12y = 568 0.12y = 568 - 160 0.12y = 408 0.12y / 0.12 = 408 / 0.12 y = 3400 so Geoff should invest $3400 at 12% and $1600 at 10% $3400 at 12% = $408 $1600 at 10% = $160 408 + 160 = $568 earned in a year dont forget to click good answer :)

OpenStudy (ucreategames):

thnx

OpenStudy (lilg132):

your welcome